OpenStudy (anonymous):

find the solution to the equation: y/4-1/4=2/3-y multiply by 12

OpenStudy (anonymous):

is it y/4 - 1/4 = 12(2/3 - y) ?

OpenStudy (anonymous):

\[12*(\frac{y}{4}-\frac{1}{4} = \frac{2}{3} - y)\]is equivalent to \[3y-3=8-12y\]\[15y=11\]\[y=\frac{11}{15}\]

OpenStudy (anonymous):

Question is unclear.